VI. Find the maximum speed of a particle that undergoes SHM with a period of 

             0.50 s and an amplitude of 20 mm. 

 

    


1
Expert's answer
2022-05-30T15:18:09-0400
"v_{\max}=A\omega=A*\frac{2\pi}{T}\\
=20\:\rm mm*\frac{2\pi}{0.50\:\rm s}=250\:\rm mm/s=0.25\: m/s"